News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 2 February 2001 8.30am EST

During the last week, Chandra completed the observing schedule as planned.

A software patch was prepared and uplinked to the spacecraft on 2/1/01 to recover data from the Solid State Recorder following a missed pass and procedural problem. The potential for data loss was recognized immediately on 1/30/01 and recording swapped from SSR-B to SSR-A. The patch to modify the values of the SSR-B pointers allowed recovery of all data from the recorder without loss or operational impact.

The planned observations for the coming week are shown below and will incorporate a real-time pass on 8 Feb to uplink the short eclipse patch in support of the start of the Spring 2001 eclipse season on 11 Feb.

------------------------------------------
Vela X-1                ACIS-S/HETG  Feb 5  
Radiation Belts
NGC 4051                ACIS-S       Feb 6  
A2256                   ACIS-S
B2 0326+39              ACIS-S
IC 1727                 ACIS-S       Feb 7  
Lambda Orionis          ACIS-I
Vela X-1                ACIS-S/HETG
Iota Orionis            ACIS-S/HETG
Radiation Belts                      Feb 8  
Iota Orionis            ACIS-S/HETG
HD 206267               ACIS-S/HETG
H 1821+643              ACIS-S/HETG  Feb 9  
OAO 1657-415            ACIS-S/HETG  Feb 10 
Radiation Belts
Capella                 ACIS-S/HETG  Feb 11 
Vela X-1                ACIS-S/HETG

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports